ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

3. Number Theory and Arithmetic Algebraic Geometry

1991 MS Classification: 3, 11, 14 Shakunle, Lere O., The Matran School of Mathematics, Berlin, Germany, and Bulckens, Anne, Deakin University, Woolstores, Australia The Divided Line and Logic Numbers The problem people outside the mathematics community have with mathematics is that they just could not see how it relates to their lives. They marvel at the wonders of science. And when told that until a science can relate its foundations to mathematics that science remains stillborn, they wonder why the music of mathematics sounds so strange and its song so remote. To show that mathematics is the foundation of knowledge is not just a catch phrase, this paper reveals how the fundamental entities of being are rooted in mathematics. It gives us a line, The Divided Line (1,2) and a number, the Logic Number (3) that lead us back to the lost paradise of mathematics where mathematics is life. References:

(1) Plato: “Four Stages of Cognition” in “The Republic of Plato”, Tr. by Francis Cornford, Oxford University Press, Oxford 1941

(2) Bulckens, Anne. : “Did Plato Ever Refer to a Section in Extreme and Mean Ratio in his Writings?” [Part I] Transfigural Mathematics, Vol. 3. No.2. 1997, pp 27 – 31.

(3) Shakunle, Lere O. : Introduction to Logic Numbers, Vol. 1. [The Matran Schools Publication, Berlin 1994]